OpenStudy (anonymous):

the derivative of y=2x^2

OpenStudy (turingtest):

use\[\frac d{dx}cx^n=c\left(\frac d{dx}x^n\right)=cnx^{n-1}\]Aand please post this in the mathematics section next time thanks